OpenStudy (anonymous):

you have a small business that shreds paper. dom can shred a truck load in 4 hrs. jim can do it in 2hrs. if dom and jim work together then they should be able to finish the job in _________ hrs.

OpenStudy (anonymous):

in 4/3 hours or 1 hr 20 min

OpenStudy (anonymous):

how do you get it??

OpenStudy (anonymous):

i think it is 1hrs and 20min

OpenStudy (anonymous):

but how do you get it explain the steps

OpenStudy (anonymous):

dom can shred a truck load in 4 hrs. jim can do it in 2hrs. See since Dom can shred a whole truck in 4 hours, he can shred 1/4 in 1 hour similarly since Jim can shred a whole truck in 2 hours, he can shred 1/2 in 1 hour So together in 1 hour they can shred = 1/4 + 1/2 = 3/4 so to shred the whole truck together, they need 4/3 hours which is equal to 1hr20min